A body of mass 10kg on a smooth inclined plane of 30
1
Expert's answer
2021-08-02T08:40:13-0400
"ma=mg\sin{30}\\a=g\sin{30}\\a=9.8\sin{30}=4.9\frac{m}{s^2}"